Can you put this in standard form? slope= 5/6;(22,12)
Answer by nerdybill(7384)

Can you put this in standard form? slope= 5/6;(22,12)
.
plug given data into "point-slope" form:
y - y1 = m(x - x1)
y - 12 = (5/6)(x - 22)
y - 12 = (5/6)x - (5/6)22
y - 12 = (5/6)x - (5/3)11
y - 12 = (5/6)x - 55/3
y = (5/6)x - 55/3 + 12
y = (5/6)x - 55/3 + 36/3
y = (5/6)x - 19/3 ("slope-intercept" form)
-(5/6)x + y = - 19/3
-(5/6)x + y = -19/3 ("standard" form)
OR
-5x/6 + y = -19/3 ("standard" form)